What is the the exact value of sec 17π/4 using periodic properties?
Answer by lwsshak3(11628)

What is the the exact value of sec 17π/4 using periodic properties?
**
let us work with the more familiar function, the cosine.
17π/4=16π/4+π/4
16π/4=4π or= two periods
moving an additional π/4 makes cos π/4=√2/2 or sec π/4=√2
ans:
sec 17π/4= √2